to find no total no. of atom, if atom present on corner ,body centre, face centre , edge centre and two atoms on each body diagonal

Grade:

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

BODY CENTRED CUBIC (BCC)

 

BCC has 8 atoms at the corners and one atom, within the body. Each corner atom makes 1/8 contribution and the contribution of atom within the body = 1

 

 No of atoms present in bcc = 1/8 × 8 (at corner) + 1(at the body centre)

 

                                           = 1+1 =2

 

FACE CENTRED CUBIC (FCC)

 

Fcc has 8 atoms at the corners and 6 atoms on the faces (one on each face)

 

Contribution by atoms at the corners = 1/8 × = 1

 

Contribution by atom on the face = 1/2 × 6 = 3

 

 Number of atoms present in fcc unit cell = 1+3 = 4

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi